# Efficient Parameter Optimisation for Quantum Kernel Alignment: A Sub-sampling Approach in Variational Training

### v1

@@ Appendix D Numerical Results
Table 9: Results for GD Optimizer on the COSET dataset in section 2.1.2. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze.
Ansatz k s ROC AUC F1 Queries Speed-up
HE 96 1 1.0000 1.0000 5796768.00 1.00
HE 32 1 1.0000 1.0000 1564640.00 3.70
HE 32 4 1.0000 1.0000 6262752.00 0.93
HE 32 8 1.0000 1.0000 12525536.00 0.46
HE 16 1 1.0000 1.0000 403184.00 14.38
HE 16 4 1.0000 1.0000 1842160.00 3.15
HE 16 8 1.0000 1.0000 3684336.00 1.57
HE 8 1 0.4902 0.5795 4408.00 1315.06
HE 8 4 1.0000 1.0000 425976.00 13.61
HE 8 8 0.7451 0.6850 921080.00 6.29
HE 4 1 0.5804 0.5795 6036.00 960.37
HE 4 4 1.0000 0.9688 75004.00 77.29
HE 4 8 1.0000 1.0000 104700.00 55.37
RA 96 1 1.0000 1.0000 5796768.00 1.00
RA 32 1 1.0000 1.0000 774624.00 7.48
RA 32 4 1.0000 1.0000 6262752.00 0.93
RA 32 8 1.0000 1.0000 12525536.00 0.46
RA 16 1 0.9529 0.8750 311664.00 18.60
RA 16 4 1.0000 1.0000 1842160.00 3.15
RA 16 8 1.0000 1.0000 3684336.00 1.57
RA 8 1 0.7843 0.6161 28792.00 201.33
RA 8 4 0.9647 0.9060 118136.00 49.07
RA 8 8 0.7569 0.7500 921336.00 6.29
RA 4 1 1.0000 0.8740 22220.00 260.88
RA 4 4 1.0000 0.9060 107164.00 54.09
RA 4 8 0.7020 0.6161 198268.00 29.24
@@
Table 10: Reference Values for AF vs D. The best Classical Kernel Represents the result of grid search as described in Appendix A. Classical NN Reference is the result from Pati et al. [2022]
ROC F1
Classical NN Reference 0.818 0.777
Best Classical Kernel 0.802 0.77
Quantum Fidelity Kernel 0.764 0.741
@@
Table 11: Results for AF vs D dataset in section 2.1.3. Grouped by optimiser (ADAM, GD and SPSA) $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze.
Optimizer Ansatz k s ROC AUC F1 Queries Speed-up
SPSA HE 1566 1 0.793 0.778 2.01 x $10^{8}$ 1.00
SPSA HE 16 1 0.786 0.778 8.10 x $10^{5}$ 248.02
SPSA HE 16 4 0.793 0.778 0.81 x $10^{6}$ 61.61
SPSA HE 8 1 0.774 0.774 4.06 x $10^{5}$ 494.25
SPSA HE 8 4 0.788 0.781 0.41 x $10^{6}$ 127.57
SPSA RA 1566 1 0.745 0.774 1.85 x $10^{8}$ 1.00
SPSA RA 16 1 0.779 0.778 6.37 x $10^{5}$ 291.11
SPSA RA 16 4 0.789 0.763 0.64 x $10^{6}$ 67.45
SPSA RA 8 1 0.775 0.774 3.18 x $10^{5}$ 583.09
SPSA RA 8 4 0.773 0.763 0.32 x $10^{6}$ 138.00
ADAM HE 1566 1 0.791 0.782 2.02 x $10^{9}$ 1.00
ADAM HE 16 1 0.777 0.774 1.10 x $10^{7}$ 182.47
ADAM HE 16 4 0.787 0.774 1.10 x $10^{7}$ 45.65
ADAM HE 8 1 0.789 0.782 5.58 x $10^{6}$ 361.05
ADAM HE 8 4 0.784 0.767 0.56 x $10^{7}$ 89.30
ADAM RA 1566 1 0.774 0.778 1.30 x $10^{9}$ 1.00
ADAM RA 16 1 0.791 0.775 4.85 x $10^{6}$ 268.40
ADAM RA 16 4 0.775 0.774 0.48 x $10^{7}$ 62.98
ADAM RA 8 1 0.774 0.778 2.24 x $10^{6}$ 580.27
ADAM RA 8 4 0.770 0.762 2.24 x $10^{6}$ 146.58
GD HE 1566 1 0.775 0.778 2.02 x $10^{9}$ 1.00
GD HE 16 1 0.789 0.776 1.13 x $10^{7}$ 178.41
GD HE 16 4 0.782 0.782 1.13 x $10^{7}$ 42.28
GD HE 8 1 0.751 0.782 5.73 x $10^{6}$ 351.83
GD HE 8 4 0.775 0.782 0.57 x $10^{7}$ 88.20
GD RA 1566 1 0.783 0.778 3.75 x $10^{8}$ 1.00
GD RA 16 1 0.791 0.782 4.68 x $10^{6}$ 80.06
GD RA 16 4 0.783 0.770 0.47 x $10^{7}$ 19.76
GD RA 8 1 0.767 0.772 2.51 x $10^{6}$ 149.21
GD RA 8 4 0.795 0.778 2.51 x $10^{6}$ 37.56
@@
Table 12: Reference Values for NBU vs AFD. The best Classical Kernel Represents result of grid search as described in Appendix A. Classical NN Reference is the result from Pati et al. [2022]
ROC F1
Classical NN Reference 0.868 0.828
Best Classical Kernel 0.859 0.809
Quantum Fidelity Kernel 0.850 0.807
@@
Table 13: Results for NBU vs AFD dataset in section 2.1.3. Grouped by optimiser (ADAM, GD and SPSA) $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze.
Optimizer Ansatz k s ROC AUC F1 Queries Speed-up
SPSA HE 2796 1 0.844 0.822 7.93 x $10^{8}$ 1.00
SPSA HE 16 1 0.834 0.826 1.13 x $10^{6}$ 702.02
SPSA HE 16 4 0.843 0.828 4.03 x $10^{6}$ 174.29
SPSA HE 8 1 0.833 0.822 5.74 x $10^{5}$ 1382.46
SPSA HE 8 4 0.836 0.828 0.57 x $10^{6}$ 350.59
SPSA RA 2796 1 0.830 0.820 7.14 x $10^{8}$ 1.00
SPSA RA 16 1 0.837 0.819 9.03 x $10^{5}$ 790.26
SPSA RA 16 4 0.833 0.820 0.90 x $10^{6}$ 194.75
SPSA RA 8 1 0.841 0.822 4.48 x $10^{5}$ 1593.03
SPSA RA 8 4 0.831 0.821 0.45 x $10^{6}$ 370.98
ADAM HE 2796 1 0.834 0.826 2.82 x $10^{9}$ 1.00
ADAM HE 16 1 0.833 0.824 1.54 x $10^{7}$ 182.64
ADAM HE 16 4 0.850 0.823 1.54 x $10^{7}$ 43.07
ADAM HE 8 1 0.826 0.824 7.71 x $10^{6}$ 365.04
ADAM HE 8 4 0.840 0.824 0.77 x $10^{7}$ 93.90
ADAM RA 2796 1 0.830 0.824 2.82 x $10^{9}$ 1.00
ADAM RA 16 1 0.841 0.821 6.27 x $10^{6}$ 449.13
ADAM RA 16 4 0.833 0.819 0.63 x $10^{7}$ 111.20
ADAM RA 8 1 0.848 0.821 3.14 x $10^{6}$ 895.85
ADAM RA 8 4 0.830 0.822 0.31 x $10^{7}$ 224.12
GD HE 2796 1 0.822 0.823 2.82 x $10^{9}$ 1.00
GD HE 16 1 0.836 0.826 1.58 x $10^{7}$ 178.66
GD HE 16 4 0.837 0.828 1.58 x $10^{7}$ 42.07
GD HE 8 1 0.835 0.828 7.92 x $10^{6}$ 355.37
GD HE 8 4 0.839 0.821 0.79 x $10^{7}$ 90.22
GD RA 2796 1 0.833 0.822 2.82 x $10^{9}$ 1.00
GD RA 16 1 0.840 0.822 6.49 x $10^{6}$ 433.50
GD RA 16 4 0.835 0.819 0.65 x $10^{7}$ 111.94
GD RA 8 1 0.843 0.824 3.26 x $10^{6}$ 862.77
GD RA 8 4 0.834 0.836 0.33 x $10^{7}$ 207.59

### v2

@@ Appendix E Numerical Results
Table 8: Results for GD Optimizer on the COSET dataset in section 2.1.2. $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze. Full kernels ($k=m$) are highlighted in bold.
[…]
0.4! Ansatz k s ROC AUC F1 Queries Speed-up HE 96 1 1 1 5796768 1 HE 32 1 1 1 1564640 3.7 HE 32 4 1 1 6262752 0.93 HE 32 8 1 1 12525536 0.46 HE 16 1 1 1 403184 14.38 HE 16 4 1 1 1842160 3.15 HE 16 8 1 1 3684336 1.57 HE 8 1 0.4902 0.5795 4408 1315.06 HE 8 4 1 1 425976 13.61 HE 8 8 0.7451 0.685 921080 6.29 HE 4 1 0.5804 0.5795 6036 960.37 HE 4 4 1 0.9688 75004 77.29 HE 4 8 1 1 104700 55.37 RA 96 1 1 1 5796768 1 RA 32 1 1 1 774624 7.48 RA 32 4 1 1 6262752 0.93 RA 32 8 1 1 12525536 0.46 RA 16 1 0.9529 0.875 311664 18.6 RA 16 4 1 1 1842160 3.15 RA 16 8 1 1 3684336 1.57 RA 8 1 0.7843 0.6161 28792 201.33 RA 8 4 0.9647 0.906 118136 49.07 RA 8 8 0.7569 0.75 921336 6.29 RA 4 1 1 0.874 22220 260.88 RA 4 4 1 0.906 107164 54.09 RA 4 8 0.702 0.6161 198268 29.24
@@
Table 9: Reference Values for AF vs D. The best Classical Kernel Represents the result of grid search as described in Appendix A. Classical NN Reference is the result from [27]
[…]
0.35! ROC F1 Classical NN Reference 0.818 0.777 Best Classical Kernel 0.802 0.77 Quantum Fidelity Kernel 0.764 0.741
@@
Table 10: Results for AF vs D dataset in section 2.1.3. Grouped by optimiser (ADAM, GD and SPSA) $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze. Full kernels ($k=m$) are highlighted in bold.
[…]
0.4! Optimizer Ansatz k s ROC AUC F1 Queries Speed-up SPSA HE 1566 1 0.793 0.778 2.01 x $\mathbf{10^{8}}$ 1 SPSA HE 16 1 0.786 0.778 8.10 x $10^{5}$ 248.02 SPSA HE 16 4 0.793 0.778 0.81 x $10^{6}$ 61.61 SPSA HE 8 1 0.774 0.774 4.06 x $10^{5}$ 494.25 SPSA HE 8 4 0.788 0.781 0.41 x $10^{6}$ 127.57 SPSA RA 1566 1 0.745 0.774 1.85 x $\mathbf{10^{8}}$ 1 SPSA RA 16 1 0.779 0.778 6.37 x $10^{5}$ 291.11 SPSA RA 16 4 0.789 0.763 0.64 x $10^{6}$ 67.45 SPSA RA 8 1 0.775 0.774 3.18 x $10^{5}$ 583.09 SPSA RA 8 4 0.773 0.763 0.32 x $10^{6}$ 138 ADAM HE 1566 1 0.791 0.782 2.02 x $\mathbf{10^{9}}$ 1 ADAM HE 16 1 0.777 0.774 1.1 x $10^{7}$ 182.47 ADAM HE 16 4 0.787 0.774 1.1 x $10^{7}$ 45.65 ADAM HE 8 1 0.789 0.782 5.58 x $10^{6}$ 361.05 ADAM HE 8 4 0.784 0.767 0.56 x $10^{7}$ 89.30 ADAM RA 1566 1 0.774 0.778 1.30 x $\mathbf{10^{9}}$ 1 ADAM RA 16 1 0.791 0.775 4.85 x $10^{6}$ 268.40 ADAM RA 16 4 0.775 0.774 0.48 x $10^{7}$ 62.98 ADAM RA 8 1 0.774 0.778 2.24 x $10^{6}$ 580.27 ADAM RA 8 4 0.770 0.762 2.24 x $10^{6}$ 146.58 GD HE 1566 1 0.775 0.778 2.02 x $\mathbf{10^{9}}$ 1 GD HE 16 1 0.789 0.776 1.13 x $10^{7}$ 178.41 GD HE 16 4 0.782 0.782 1.13 x $10^{7}$ 42.28 GD HE 8 1 0.751 0.782 5.73 x $10^{6}$ 351.83 GD HE 8 4 0.775 0.782 0.57 x $10^{7}$ 88.20 GD RA 1566 1 0.783 0.778 3.75 x $\mathbf{10^{8}}$ 1 GD RA 16 1 0.791 0.782 4.68 x $10^{6}$ 80.06 GD RA 16 4 0.783 0.77 0.47 x $10^{7}$ 19.76 GD RA 8 1 0.767 0.772 2.51 x $10^{6}$ 149.21 GD RA 8 4 0.795 0.778 2.51 x $10^{6}$ 37.56
@@
Table 11: Reference Values for NBU vs AFD. The best Classical Kernel Represents result of grid search as described in Appendix A. Classical NN Reference is the result from [27]
[…]
0.35! ROC F1 Classical NN Reference 0.868 0.828 Best Classical Kernel 0.859 0.809 Quantum Fidelity Kernel 0.850 0.807
@@
Table 12: Results for NBU vs AFD dataset in section 2.1.3. Grouped by optimiser (ADAM, GD and SPSA) $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze. Full kernels ($k=m$) are highlighted in bold.
[…]
0.5! Optimizer Ansatz k s ROC AUC F1 Queries Speed-up SPSA HE 2796 1 0.844 0.822 7.93 x $\mathbf{10^{8}}$ 1 SPSA HE 16 1 0.834 0.826 1.13 x $10^{6}$ 702.02 SPSA HE 16 4 0.843 0.828 4.03 x $10^{6}$ 174.29 SPSA HE 8 1 0.833 0.822 5.74 x $10^{5}$ 1382.46 SPSA HE 8 4 0.836 0.828 0.57 x $10^{6}$ 350.59 SPSA RA 2796 1 0.83 0.820 7.14 x $\mathbf{10^{8}}$ 1 SPSA RA 16 1 0.837 0.819 9.03 x $10^{5}$ 790.26 SPSA RA 16 4 0.833 0.82 0.90 x $10^{6}$ 194.75 SPSA RA 8 1 0.841 0.822 4.48 x $10^{5}$ 1593.03 SPSA RA 8 4 0.831 0.821 0.45 x $10^{6}$ 370.98 ADAM HE 2796 1 0.834 0.826 2.82 x $\mathbf{10^{9}}$ 1 ADAM HE 16 1 0.833 0.824 1.54 x $10^{7}$ 182.64 ADAM HE 16 4 0.85 0.823 1.54 x $10^{7}$ 43.07 ADAM HE 8 1 0.826 0.824 7.71 x $10^{6}$ 365.04 ADAM HE 8 4 0.840 0.824 0.77 x $10^{7}$ 93.90 ADAM RA 2796 1 0.83 0.824 2.82 x $\mathbf{10^{9}}$ 1 ADAM RA 16 1 0.841 0.821 6.27 x $10^{6}$ 449.13 ADAM RA 16 4 0.833 0.819 0.63 x $10^{7}$ 111.20 ADAM RA 8 1 0.848 0.821 3.14 x $10^{6}$ 895.85 ADAM RA 8 4 0.830 0.822 0.31 x $10^{7}$ 224.12 GD HE 2796 1 0.822 0.823 2.82 x $\mathbf{10^{9}}$ 1 GD HE 16 1 0.836 0.826 1.58 x $10^{7}$ 178.66 GD HE 16 4 0.837 0.828 1.58 x $10^{7}$ 42.07 GD HE 8 1 0.835 0.828 7.92 x $10^{6}$ 355.37 GD HE 8 4 0.839 0.821 0.79 x $10^{7}$ 90.22 GD RA 2796 1 0.833 0.822 2.82 x $\mathbf{10^{9}}$ 1 GD RA 16 1 0.84 0.822 6.49 x $10^{6}$ 433.50 GD RA 16 4 0.835 0.819 0.65 x $10^{7}$ 111.94 GD RA 8 1 0.843 0.824 3.26 x $10^{6}$ 862.77 GD RA 8 4 0.834 0.836 0.33 x $10^{7}$ 207.59